OpenStudy (anonymous):

the sum of three consecutive number is 72. What are the smallest of these numbers?

OpenStudy (camper4834):

the smallest is when all the numbers are the same

OpenStudy (camper4834):

if they are all the same you could say this number is x

OpenStudy (camper4834):

oh consecutive

OpenStudy (camper4834):

well lets keep going

OpenStudy (camper4834):

this number will be x

OpenStudy (camper4834):

that means 3x=72

OpenStudy (anonymous):

Let x = first number, x + 1 = second number, x + 2 = third number. x + (x + 1) + (x + 2) = 72 x + x + 1 + x + 2 = 72 3x + 3 = 72 3x + 3 - 3 = 72 - 3 = 69 3x/3 = 69/3 x = 23 The smallest of the three numbers is 23. Verify 23 + 24 + 25 = 60 + 12 = 72